OpenStudy (anonymous):

Approximate rad 55 to 3 decimal places do not use calculator use only binomial expansion Theorem i got up to (49+6)^(1/2)

OpenStudy (anonymous):

(49+6)^(1/2) = n C o W^(1/2) plug it in to (W+U)^n which is nCoW^n + nC(1) W^(n-1)U ...... +nCn U^n will this work?

OpenStudy (dumbcow):

this is what you want http://en.wikipedia.org/wiki/Binomial_theorem#Generalisations

OpenStudy (dumbcow):

\[(49+6)^{1/2} = \sqrt{49} + \frac{1}{2}\frac{6}{\sqrt{49}}-\frac{1}{8} \frac{6^2}{(\sqrt{49})^3} ...\]

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw